A Sporting goods store a pair of skates for 102. The skates cost the store 68. What percent of increase was this of the original cost?
Choices 50 percent or "33 1/3" percent

Answer by drk(1908)

percentage increase is (end-start)/start.
% increase = (102-68)/68 = 34/68 = 50%